### v1

# Cyber-physical WebAssembly:

Abstract
The rapid expansion of Internet of Things (IoT), edge, and embedded devices in the past decade has introduced numerous challenges in terms of security and configuration management. Simultaneously, advances in cloud-native development practices have greatly enhanced the development experience and facilitated quicker updates, thereby enhancing application security. However, applying these advances to IoT, edge, and embedded devices remains a complex task, primarily due to the heterogeneous environments and the need to support devices with extended lifespans. WebAssembly and the WebAssembly System Interface (WASI) has emerged as a promising technology to bridge this gap. As WebAssembly becomes more popular on IoT, edge, and embedded devices, there is a growing demand for hardware interface support in WebAssembly programs. This work presents WASI proposals and proof-of-concept implementations to enable hardware interaction with I2C and USB, which are two commonly used protocols in IoT, directly from WebAssembly applications. This is achieved by running the device drivers within WebAssembly as well. A thorough evaluation of the proof of concepts shows that WASI-USB introduces a minimal overhead of at most 8% compared to native operating system USB APIs. However, the results show that runtime initialization overhead can be significant in low-latency applications.
Index Terms:
WebAssembly, WASI, Edge, Embedded, IoT, Cyber-physical, Hardware Abstraction Layer
I Introduction
The Internet of Things (IoT) refers to the network of devices and electronics embedded with various sensors, software, and other technologies to gather and exchange information with other devices and systems over the internet [1]. During the past decade, IoT has experienced remarkable growth in a variety of industries and has become an integral part of our daily lives with more than 17 billion devices connected in 2024 [2].
A large amount of IoT devices are integral to critical infrastructure [3] and operate on the edge of the internet, where the management of their software and firmware poses significant challenges due to their inherent complexity [4]. In accordance with the Cyber Resilience Act [5], manufacturers need to ensure, among various factors, that for the duration of the support period of the product, vulnerabilities are handled effectively and that security updates should be available to users for at least the time the product is expected to be in use. Moreover, the automotive industry is increasingly adopting the practice of wirelessly distributing software updates to vehicles, known as over-the-air updates, where security is of critical importance [6, 7]. In parallel, Industrial Internet of Things systems often feature devices with operational lifespans of more than 30 years [8, 9]. To ensure forward compatibility of these systems, they must be able to integrate with newer hardware components, such as modern and updated sensors and displays, since the original hardware usually already reached end of life when they need to be replaced. This highlights the necessity for modularity at the driver level. Additionally, these systems are frequently tightly integrated with devices that have limited hardware capabilities, which augments the requirements with the demand for minimal overhead.
Although cloud-native development practices have introduced benefits such as isolation, componentization, containerization, and modern deployment and orchestration platforms, these advances are difficult to apply to edge environments and even more so for IoT devices, embedded systems, and microcontrollers [10]. Ensuring drivers are up-to-date and secure is crucial for supply chain security, as they frequently have elevated system privileges or direct access to hardware. W3C [11] WebAssembly (Wasm) [12] and the WebAssembly System Interface (WASI) [13] emerges as a promising solution to shrink the trusted computing base and further enhance embedded development and maintainability by providing a trusted runtime, modern development methodologies, toolchains and software development kits. Wasm was originally created to execute binary code within the browser to improve the performance of web applications [14], and is now actively used outside the browser via WASI [15]. WASI facilitates operating system communication within Wasm by providing a set of portable Application Programming Interfaces (APIs). However, applications on embedded and IoT devices require access to hardware interfaces to enable communication with the system in which they operate. Wasm does not currently meet the demands of cyber-physical systems, particularly in accessing connected hardware, and requires standardized WASI interfaces to be implemented in Wasm runtimes to maintain its portability and enable communication with hardware across diverse devices.
In this work, we propose standardized WASI interfaces and proof-of-concept implementations for connecting hardware using I2C and USB, along with security mediation and isolation, on low-powered embedded and IoT devices running a Wasm runtime with pluggable drivers. Given that native drivers are inherently insecure due to the fact that they involve raw memory manipulation, our solution aims to reduce the amount of components that need to be trusted, thus enhancing supply chain security. Furthermore, our proposal emphasizes the importance of forward compatible devices and modularity at the driver level, ensuring that the system remains adaptable and secure in the face of evolving technological landscapes. Specifically, this paper aims to answer the following questions:
RQ 1: How to enable WebAssembly applications outside the browser to interact with hardware interfaces such as I2C and USB?
RQ 2: How to apply access control security measures to ensure authorized access to hardware interfaces?
RQ 3: How to standardize cyber-physical WebAssembly interfaces to ensure future compatibility?
RQ 4: How does interacting with hardware interfaces through WebAssembly impact performance compared to using native solutions?
The remainder of this paper is structured as follows: Section II investigates the state of the art concerning the use of WebAssembly on IoT and edge devices. Section III details our proof-of-concept architecture for supporting I2C and USB hardware interfaces within WebAssembly while Section IV discusses the standardization efforts for the WASI proposals, highlighting key decisions made during their development. Section V illustrates the feasibility of the WASI proposals through the development of multiple proof-of-concept implementations and Section VI presents the evaluation results of these implementations. Finally, Section VII concludes the paper.
II Related Work
WebAssembly has emerged as a promising technology for enabling high-performance, cross-platform applications, particularly in the context of embedded systems and IoT devices. Several studies have explored the integration of Wasm with various hardware and software environments to enhance the capabilities and security of these systems, which are discussed below.
Singh et al. [16] developed the WARDuino virtual machine (VM), which exposes Arduino APIs for GPIO (General-Purpose Input/Output), SPI (Serial Peripheral Interface), ADC/DAC (Analog to Digital and Digital to Analog Converter), and PWM (Pulse Width Modulation) to the Wasm runtime. WARDuino only facilitates the use of Wasm in, as its name suggests, Arduino-based projects. This poses notable limitations in terms of portability. This also means that the hosting device must provide all the necessary APIs and that there is no hardware support beyond the Arduino ecosystem.
Aerogel [17] is a lightweight access control framework that addresses security gaps between the bare-metal IoT devices and the Wasm execution environment concerning access control for sensors, actuators, processor energy usage, and memory usage. Aerogel enhances the security and provides resource management in multi-tenant Wasm runtime environments, where each Wasm application is considered a tenant. Aerogel allows sensor and actuator interaction by exporting high-level sensing and actuation functions to the WebAssembly Micro Runtime (WAMR) [18]. A user-defined access control specification sheet allows one to set restrictions on specific I/Os, limitations on the number of accesses within a specified time frame, an upper bound on memory and CPU usage, and the maximum number of applications having shared access to a specific hardware component. Access to the requested resource will only be granted if all access control conditions are met.
WiProg [19] is a framework designed to simplify integrated IoT programming and application development for sensor-to-edge systems by employing Wasm, with a focus on application offloading. It allows for a monolithic application development approach which allows WiProg to perform offloading of the application based on compiler annotations and offloading policies. WiProg provides hardware-agnostic APIs to access peripherals on IoT devices, but these are limited to analog I/O, digital I/O, and Universal Asynchronous Receiver-Transmitter (UART) APIs, thus lacking support for other important protocols like I2C.
Li et al. [20] proposed WAIT, a lightweight Wasm runtime for resource-constrained IoT devices for device-cloud integrated applications. It enables Wasm ahead-of-time (AOT) compilation targeting resource-constrained devices. When a device running WAIT receives a Wasm module binary, it performs on-device AOT compilation, checks sandboxing guarantees, and finally executes the module. To cope with the limited resources, the compile-time memory footprint is reduced by streamed look-back compilation. To access peripherals and I/O, native APIs are provided by the WAIT runtime and are exposed via the import mechanism of Wasm. The AOT compiler then performs binary rewriting techniques to replace the instructions that call the imported function with assembly code that directly accesses the peripheral, reducing the overhead associated with calling an external function.
Finally, Wasmachine [21] is a bare-metal operating system designed for embedded devices. It operates without protection rings by running everything in kernel mode, allowing for zero-cost system calls. The Wasmachine kernel exports system calls in WASI prototypes and runs each Wasm application as a kernel thread, allowing the application to invoke system calls as normal function calls without incurring additional context switching costs. Wasmachine allows access to hardware by using WASI, where access permissions are configured on a per-application basis and none are granted by default. Despite its innovative approach, the system has certain limitations, including limited support of CPU architectures and the reliance on a Unix-like monolithic kernel architecture, which can limit the feasibility on certain IoT or edge devices.
III Architecture
Fig. 1 shows a diagram of the proof-of-concept architecture designed by the authors to enable I2C and USB hardware interface support in Wasm applications. Both the application and the peripheral device drivers run as WebAssembly components. These connect to the runtime via standardized generic I2C and USB WebAssembly System Interfaces. The runtime itself has a number of host components, one for each interface, that implement the interfaces and provide access-control list (ACL) and capability-based security. The host component itself communicates with the host operating system (OS) and uses OS-specific APIs for hardware access. The remainder of this section examines important concepts in more detail.
Figure 1: Schematic representation of the proof-of-concept architecture that enables I2C and USB hardware interface support in Wasm applications. The application and device drivers run as Wasm guest components and connect to the host components via WASI. The host components provide ACL and capability-based security and use underlying OS-specific APIs to implement the interfaces. This figure contains resources from Flaticon.com.
III-A Interface Worlds
An interface world describes the system calls that are supported by the runtime using the Component Model’s Interface Description Language (IDL) called Wasm Interface Type (WIT). For example, the USB world contains a WIT description of all the system calls that Wasm components can use to communicate with USB devices.
This world is a high-level interface contract between the runtime and the guest component and can be used by tools such as wit-bindgen [22] to automatically generate bindings for calling these APIs from various languages. The source code of the guest components uses these generated bindings to implement the world and to call the imported APIs, allowing them to interact with hardware devices supported by the interface. The runtime host component can use these bindings to implement and export the functions that make up the API.
III-B Host Component
The host component is the host-side implementation of the system call API. The host component is embedded in the runtime and runs as native code directly on the underlying machine, as it requires access to native APIs of the host operating system to communicate with hardware devices. As such, the host component is highly dependent on the specific host OS interfaces. Because it is embedded in the runtime, the host component also depends on the specific implementation details of the runtime. Therefore, the implementation of the host component is unique for every combination of runtime and underlying (operating) system. Any new system or runtime that wishes to support certain worlds, will require a modified implementation of the host component. However, this burden can be lessened by using higher-level cross-platform libraries to implement the host component, rather than depending directly on the operating system through system calls. This way, one implementation of the host component can be compiled for multiple operating systems. Depending on the specific implementation, it may also be possible to target multiple processor architectures running one of the supported operating systems using a single implementation. Unfortunately, using a higher-level library is not always feasible, especially on certain low-powered embedded devices due to device-specific (operating system) limitations, requiring a custom host component implementation for those devices.
III-C Guest Components
Guest components contain the application logic and device drivers or a hardware abstraction layer (HAL). These can be made up of multiple separate components, or the application logic can embed the driver and HAL logic inside of it. A guest component specifies which interfaces it requires through imports and the interfaces it provides through exports using a WIT world description. This informs the runtime about which host component interfaces to connect to the guest component, allowing the guest component to interact with hardware devices through host component function calls. Guest components are contained within a sandbox by the runtime, ensuring they remain isolated and cannot access system resources without authorization.
III-D Runtime
The Wasm runtime resolves the imports and exports for all the components and links them together. To successfully resolve imports and exports, the runtime must support the WIT world that a component implements. For example, this can be achieved by having a host component that provides implementations for the imported functions required by the guest component. After linking, the guest component has access to the imported functions from the world it implements. The runtime also has access to the functions that the world exports. To begin execution, the runtime must call one of those exported functions. Since interface worlds are not concerned with how a component starts execution, this is delegated to another world. More specifically, the runtime supports the wasi-cli [23] world to specify the entry-point of the guest component. Popular Wasm runtimes include Wasmtime [24] and WAMR [18], which are both being developed by the Bytecode Alliance [25].
III-E Capability-Based Security
One of the strengths of WASI is its capability-based security mechanism. This mechanism implicitly allows granular control over what resources an application can access based on the API definition. If a resource is not provided, the module cannot interact with it, implicitly revoking the capability for that resource. For example, WASI file system APIs apply capability-based security so that applications can only access files and directories to which they have been granted access [26]. Otherwise, the resource does not exist.
Although WASI enables the implementation of capability-based security for system interfaces, it does not stipulate the level of granularity at which these security controls should be constructed. To illustrate, a USB API may employ capability-based access control at multiple levels:
  • Device level: Allow only devices with specific vendor and product ID pairs.
  • Configuration level: Allow only devices with specific configuration types. For example, allow only self-powered devices.
  • Interface level: Allow only interfaces with specific property types. For example, allow only interfaces that have the mass storage class code.
  • Endpoint level: Allow only endpoints with specific property types. For example, allow only endpoints with an IN direction that restricts the guest to receiving data only.
The capability-based security model is inherently defined by the API, which also determines the granularity of the security measures. This implies that any access control measures must define the resources to be protected, which is impractical for fine granularity and sometimes impossible when control layers are orthogonal or complementary. Consequently, we propose an extension of capability-based security with explicit access control list security.
III-F ACL-Based Security
Access control lists provide an additional layer of granular and orthogonal access control, complementing capability-based access control. Unlike capability-based access control, which requires changes to the API, ACLs can be implemented without affecting the existing API. Furthermore, ACLs offer greater flexibility and can be fine-tuned to the environment and runtime as needed. Based on these advantages, we propose that standardized interfaces use capability-based security for high-level concepts where the benefits of access control are clear and use ACL-based access control to further restrict permissions.
The implementation of ACL-based access control necessitates the incorporation of logic specific to the runtime, environment, and peripheral. Whenever the runtime creates a new module instance, an I/O context is created and linked to the corresponding execution environment. Whenever a device interaction is performed, the corresponding I/O context is retrieved from the execution environment. Depending on the type of device interaction and the I/O context’s properties, the action is permitted or denied.
The I/O context’s properties may refer to permitted interaction types, such as read or write. Additionally, specific values may be blocked, which is necessary for device interaction with arbitrary data, but, for example, prevents changes in the configuration. Furthermore, the context may specify how connections should be established with end devices, preventing communication with other peripherals sharing the same bus.
IV System Interface Standards
In order to ensure compatibility with future toolchains and platforms, it is important to standardize the cyber-physical WebAssembly interfaces. Standardization will enable the integration of support for these interfaces into runtimes, compiler toolchains, HAL libraries, and drivers.
As part of this research, the authors joined the WebAssembly System Interface Subgroup of the W3C WebAssembly Community Group [27] to initiate the standardization process [^1] https://github.com/WebAssembly/meetings/blob/main/process/phases.md for these interfaces. These efforts focused on two interfaces commonly used in IoT and embedded systems: I2C and USB. At the time of writing, the WASI-I2C [28] proposal is in phase 2 and the WASI-USB [29] proposal is in phase 1 of the standardization process. In the following, we outline both proposals, emphasizing certain key decisions that were made during the development of the proposals.
IV-A WASI-I2C
The WASI-I2C proposal is co-championed by representatives from both imec and Siemens, and is currently in the second phase of the standardization process, with ongoing efforts to fulfill the criteria to pass the vote to the third phase.
The WASI-I2C proposal contains three WIT files adhering to the component model: delay, i2c, and world. The delay and i2c files are closely aligned with the interfaces from embedded-hal [30] v1.0, which served as inspiration, although the use of the embedded-hal Rust crate is not required to implement the proposal. The world file provides a way to import all the described interfaces of the proposal at once, thus currently importing all interfaces from the delay and i2c files.
The proposal describes two handles that dynamically identify and provide access to resources, similar to a file descriptor. One handle is designated for I2C and the other for delays. These handles offer extensive access to the resource, meaning you currently cannot restrict the I2C connection to specific addresses. While a more fine-graded control model is feasible, consultations with the embedded special interest group (SIG-Embedded) [31] and industry partners have indicated that there is currently no demand for this.
Because there exist various ways to obtain an I2C handle, the i2c interface lacks an explicit constructor for the i2c resource. Instead, worlds that utilize this interface should provide their own mechanism to obtain I2C handles, typically by passing handles into exported functions as parameters or by having imported functions return them. The i2c resource includes essential methods such as read and write to interact with the I2C device.
Although delaying program execution is not inherently associated with I2C communication, certain guest implementations depend on it, which is why it is also included in the current proposal. In the future, delays could either be separated into their own proposal or merged with other embedded-related proposals to avoid the need for each delay dependent proposal to establish its own standards. The main downside of merging proposals is that it could significantly increase the workload for the proposal champions, who would need to be interested and committed to maintaining these additional communication protocols. However, having centralized proposals could streamline the process for potential implementers.
A key element of a proposal is its portability criteria. These criteria demonstrate that the champions’ specification is not overly tailored to a specific platform. Typically, there need to be at least two entirely independent implementations in separate Wasm runtimes. However, given the diversity of platforms that can interface with I2C, this approach is not feasible for the WASI-I2C proposal. The champions, along with representatives from Siemens, Intel, Aptiv, Xiaomi, Amazon, Midokura, Sony, Atym, and Bosch, have collectively established a set of portability criteria for the proposal, as outlined in Table I. To minimize the potential variety in memory size, flash size, and processor speed, reference hardware was provided for each criterion. Although the Intel x86 architecture is not explicitly mentioned, its support is implied since Wasmtime is being developed using this architecture. The NUCLEO-F412ZG was selected as it serves as Siemens’ reference board. The remaining two targets were selected to demonstrate compatibility with both a real-time operating system (RTOS) and Linux, as well as ARM and RISC-V architectures. In addition to the target requirements, it is crucial that the interface is designed in a way that minimizes its memory usage, ensuring that there is sufficient memory available for the application itself.
TABLE I: WASI-I2C Proposal Portability Criteria
Platform Architecture Reference Hardware
Linux ARM64 Raspberry Pi 3B
RTOS RISC-V ESP32-C2
RTOS ARM32 NUCLEO-F412ZG
IV-B WASI-USB
The WASI-USB proposal is championed by imec and is currently in the first phase of the standardization process.
The WASI-USB proposal contains a preliminary version of the USB API interface and is described using the WIT format. The key components of the proposal are outlined below.
The usb interface contains a usb-device and a device-handle resource. The static enumerate function of the usb-device resource allows the caller to retrieve all available USB devices on the bus, returning them as a list of usb-device instances. Moreover, there are three other methods that can be used on any of these instances. The device-descriptor and configurations methods provide additional information about a device and are always available without requiring exclusive access to the device. Conversely, the open method is only available under specific conditions. This method establishes a context for device communication. It is important to note that opening a device can fail for various reasons. For instance, some operating systems may refuse to open a device if the executable lacks the necessary permissions. These permissions are managed at the Wasm runtime level, rather than being specific to individual Wasm modules. To develop a robust and user-friendly API, functions that require an ‘open’ device are scoped to the device-handle resource, which is only obtainable when the device is opened successfully. The device-handle resource encompasses all the methods necessary for device communication, with most methods closely related to similar libusb [32] functions.
The events interface allows a guest to poll for connection event updates. This allows the guest to detect connect and disconnect events and react accordingly. The events interface includes a single update function that returns a device-connection-event variant. The returned case of device-connection-event will vary based on the specific event that has occurred, for example pending when no new event update is available. With the anticipated async support in WASI Preview 3 [33, 34], the update function can be adapted to receive event updates through an asynchronous stream rather than polling.
IV-B1 Security Model Considerations
The current WASI-USB proposal applies the capability-based security model exclusively at the device level. In this approach, only devices with specified vendor and product ID pairs are permitted. This is the most straightforward option and is similar to how WebUSB protects devices [35]. Community feedback will be essential to determine whether the capability-based security model should be extended to other levels, based on its perceived usefulness.
IV-B2 Exclusive Device Access
Multiple components may attempt to access the same USB device simultaneously. However, this should be avoided, as each component typically expects exclusive access to a device. To address this, the runtime should monitor which components are using specific devices and prevent other components from accessing a device that is already in use. This can be achieved by returning an error when a component attempts to open a device that is already in use. This approach is consistent with most operating systems, where only one program can access a USB device at any given time.
V Implementation Details
To demonstrate the feasibility of the proposed cyber-physical WebAssembly architecture, we developed several proof-of-concept implementations, each addressing different aspects of the proposed solution. Our primary objectives were to validate the proposed approach and identify potential opportunities or shortcomings in the underlying architecture and standard methodologies. In particular, this research presents three distinct proof-of-concept implementations.
  • I2C in Wasmtime: Demonstrates driving an I2C seven-segment display from WebAssembly running on a Linux host and on a Raspberry Pi Pico microcontroller, as well as reading temperature and humidity data from an HTS221 sensor on a Raspberry Pi 3B.
  • I2C in WAMR: Provides a proof of concept that can read data from an HTS221 sensor. Note that this solution does not yet use the component model or capability-based security due to the current limitations of WAMR, including the lack of support for WASI Preview 2. Instead, an ACL-based Security System was implemented.
  • USB in Wasmtime: Showcases the communication of WebAssembly applications with USB devices, including a USB flash drive using a WebAssembly-based driver and an Xbox controller using a separate driver.
The remainder of this section describes each proof-of-concept implementation.
V-A I2C in Wasmtime
The I2C in Wasmtime proof-of-concept implementation [36] is made up of several components: the host implementation, two WIT worlds and two guest implementations.
The host implementation handles the actual I2C connection with the device, whereas the guest implementations are responsible for device-specific logic, such as reading the temperature or writing the time to the display. The host and guest components interact with each other through two WIT worlds, one for the seven-segment display and the other for the HTS221 sensor. These WIT worlds import the necessary interfaces from the WASI-I2C proposal that are implemented by the host, allowing the guest to communicate with the I2C device. Additionally, these WIT worlds export device-specific functions, such as get-temperature, that are implemented and made available by the guest. Although these worlds were written for our specific proof of concepts, they aim to be as generic as possible to allow future support for other displays and sensors. For example, the display’s write function can accept any string, even though our display only has four seven-segment characters.
The guest implementations do not make use of Rust’s std library because they only need specific types that can be imported individually, which also helps minimize the size of the compiled Wasm binaries. The guest implementation for the seven-segment display directly writes the necessary bytes to control it, while the guest component for the HTS221 sensor leverages the embedded-hal compatible hts221 [37] Rust crate for simplicity. To adapt the embedded-hal API to the slightly different generated bindings, the wasi-embedded-hal [38] crate was developed.
V-B I2C in WAMR
WAMR currently does not support WASI Preview 2. Consequently, the component model and WIT cannot be used in WAMR, which means that a standardized I2C connection cannot be passed to the guest because WIT resources are not available. As a workaround, the I2C connection is kept global on the host side. For the host to be able to bind with the guest, it has to register the function that the guest imports and find the functions that the guest exports. Parameters are passed to the guest’s exported functions using a vector of WasmValue. Internally, a WasmValue is represented as an array of 32-bit unsigned integers.
The ACL security system for the I2C implementation employs a struct i2c_context to store access permissions. A reference is registered in the user_data field of the execution environment, which contains the device address, indicating that the specific module is permitted to communicate with the device. As it is necessary to reestablish the I2C connection on each I2C API call, a check for permissions can be performed. Given the frequency with which I2C functions are called in embedded applications, it is necessary to ensure that the lookup time is constant. To this end, the allowed addresses are stored in two 64-bit fields spanning the full range of 128 valid I2C addresses, thereby ensuring that the lookup time remains consistent.
V-C USB in Wasmtime
The proof-of-concept implementation [39] of USB in Wasmtime includes a host implementation along with several guest implementations, such as interfacing with a USB flash drive or an Xbox controller. The main concept of the USB implementation in Wasmtime mirrors that of the I2C implementation in Wasmtime: the host component is platform dependent and interacts with the underlying system to access the USB resource, while the guest components handle device-specific logic, such as reading the joystick positions of an Xbox controller.
To facilitate the development of the host component and enhance its platform independence, rusb [40], a Rust wrapper around libusb [32], is used. The USB flash drive guest component implements a range of commands for USB Attached SCSI devices, including basic file system operations like listing directory contents and reading files. To streamline the proof-of-concept implementation, it utilizes the fatfs [41] Rust crate for file system support. The file system support is fully integrated within the guest component itself, without relying on the host operating system’s file system capabilities. The Xbox guest component continuously monitors the current state of the controller through its USB interrupt pipe and writes button presses and joystick positions to standard output.
VI Evaluation Results
This section assesses the practicality of our proposals in real-world scenarios. Initially, each proof of concept is evaluated to ensure it meets the functional requirements, typically by manually verifying device functionality. Subsequently, various benchmarks are performed to identify any performance implications.
VI-A I2C Performance Evaluation
The I2C proof of concept for the seven-segment display is evaluated on a Raspberry Pi 4B, whereas the HTS221 sensor proof of concept is evaluated on a Raspberry Pi 3B. The display evaluation will compare the native, Wasmtime, and WAMR implementations, while the HTS221 sensor evaluation will compare the native and Wasmtime implementations. All benchmark results are available on GitHub [36].
To assess performance overhead, we conduct 100 iterations of writing a four-digit number to the display and also reading the temperature from the HTS221 sensor. To highlight the worst-case overhead, cold start execution times are measured, where each measurement includes the time required to instantiate the Wasm binary within the runtime. The mean cold start execution times are then compared across the native, Wasmtime, and WAMR implementations. Fig. 2 illustrates the mean cold start execution times for the two proof of concepts. Wasmtime shows a notable overhead relative to the native implementation, with the mean cold start execution times being 7.16 ms (+238.7%) longer for the display and 1.38 ms (+18.9%) longer for the HTS221 proof of concept. In contrast, the WAMR implementation incurs only a minimal overhead of 50 $\mu$s (+1.7%). Most of the additional time consumed by Wasmtime can be attributed to functions associated with Cranelift [42], a compiler backend utilized by Wasmtime for both just-in-time and ahead-of-time (AOT) compilation.
Figure 2: Mean cold start execution times for writing a four-digit number to the display and reading the temperature from the HTS221 sensor, comparing native, Wasmtime, and WAMR implementations.
VI-B USB Performance Evaluation
The USB proof of concept is evaluated on three systems: one with an AMD Ryzen 7 5800H and USB3.0 running x86 Linux 6.8.9, a Raspberry Pi 3B+ with USB2.0 running AArch64 Linux 6.8.0, and a third with an AMD Ryzen 7 5800X3D and USB3.0 running x86 Windows 10 22H2. All benchmark results are available on GitHub [43].
To assess the runtime size and memory usage overhead introduced by the USB host component, we compare the Wasmtime runtime both with and without the USB host component. Our measurements show that the runtime size overhead is minimal, adding only 160 KiB (+1.1%) for the x86 Linux runtime, 220 KiB (+1.7%) for the AArch64 Linux runtime, and 275 KiB (+1.9%) for the Windows runtime. The memory usage overhead is also minimal for the Linux runtimes, comprising only 459 KiB (+1.7%) for the x86 Linux runtime and 295 KiB (+1.5%) for the AArch64 Linux runtime. The Windows version of the runtime incurs a more severe memory overhead of 1427 KiB (+11.0%), which can be attributed to an entirely different backend being used by libusb on Windows.
The impact on throughput is assessed using a USB flash drive to read or write as many data blocks as quickly as possible. Fig. 3 depicts the results for a sequential read benchmark. On x86 Linux and Windows, the overhead is minimal, respectively resulting in only a 0.8 MiB/s (-0.6%) and 0.4 MiB/s (-0.2%) reduction in median throughput. The overall throughput on AArch64 Linux is notably lower, along with a 1.4 MiB/s (-4.1%) decrease in median throughput in Wasm. This can be attributed to the fact that the Raspberry Pi 3B+ only supports USB2.0, and this also implies that the overhead is larger on less powerful hardware. Despite this, the overhead remains relatively modest. Benchmarks for sequential write, random read, and random write show similar relative overheads.
Figure 3: USB flash drive sequential read speeds across various platforms for both native and Wasm guest applications show minimal overhead on high-powered machines, while lower-powered devices experience a larger overhead. The reported speeds represent the median throughput.
The latency overhead is evaluated using an Arduino Nano 33 IoT device with custom firmware that receives a message from the host device and echoes it back to the host without modifications. This allows the host to measure the round-trip time (RTT) of this message. Fig. 4 shows the observed RTTs for a bulk transfer of a 32-byte message. Due to the low priority of bulk transfers among USB transfer modes, numerous outliers are present. Some outliers have been excluded to facilitate the comparison of the median RTTs. The overhead introduced by WebAssembly is relatively modest. On x86 Linux, it results in a median RTT that is only 7 $\mu$s (+3.5%) longer than the native implementation. On x86 Windows, an even lower overhead of 2 $\mu$s (+0.8%) is observed. However, on AArch64, the overhead is more pronounced, with an increase of 45 $\mu$s (+8.0%).
Figure 4: Round-trip time for a 32-byte USB bulk transfer across different platforms, comparing native and Wasm guest applications. x86 Linux shows minimal outliers and reasonable overhead, whereas AArch64 displays greater overhead. The reported times represent the median round-trip time.
VII Conclusion
With the growing popularity of WebAssembly on embedded and IoT devices, the need for standardized hardware interface support in WebAssembly arose. This paper introduces two WebAssembly System Interface proposals, WASI-I2C and WASI-USB, enabling the use of I2C and USB hardware interfaces within WebAssembly. The cyber-physical WebAssembly proof-of-concept architecture is described in depth, including a solution for authorized hardware interface access. Hardware interface access is managed through the capability-based security mechanism of WASI, and is extended with explicit ACL-based security for more granular access control. To ensure our proposals remain compatible with future toolchains and platforms, we have initiated the standardization process in collaboration with industry partners. Development details for the standardized proposals are provided, highlighting key decisions made throughout the process. Proof-of-concept implementations showcase the feasibility of the proposed architecture and are evaluated to confirm that they meet functional requirements and to identify any performance impacts. Our evaluation results show that the overhead introduced by WASI-USB is minimal, resulting in a maximum throughput reduction of 1.4 MiB/s (-4.1%) and an increase in round-trip time of up to 45 $\mu$s (+8.0%). Our test results for WASI-I2C include the initialization of the Wasm binary within the runtime in each measurement. This reveals significantly worse overhead in Wasmtime, with the mean cold start execution time increasing by up to 7.16 ms (+238.7%) at worst and 1.38 ms (+18.9%) at best. This emphasizes that the overhead of runtime and binary initialization is significant, indicating that low-latency applications require a warm-up period.
This work has established a foundation for standardizing the I2C and USB WASI interfaces, and demonstrates the feasibility of a cyber-physical WebAssembly future on IoT and embedded devices. However, future work will be necessary to stabilize the standardization proposals, which will involve extensive testing, implementation across multiple runtimes, and gathering community feedback. Additionally, these proposals lack asynchronous support, which limits their functionality in certain environments.
Acknowledgment
We extend our heartfelt gratitude to everyone who contributed to these proposals: Valentin Olpp, Dan Gohman, Emiel Van Severen and the Bytecode Alliance & WASI subgroup of the W3C WebAssembly Community Group.
This work received funding from the Smart Networks and Services Joint Undertaking (SNS JU) under the Horizon Europe research and innovation program of the European Union under Grant Agreement number 101139067 (https://elasticproject.eu/). Expressed views and opinions are however those of the authors only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@
 Index Terms:
-WebAssembly, WASI, Edge, Embedded, IoT, Cyber-physical, Hardware Abstraction Layer
+Cyber-physical, edge, embedded, hardware abstraction layer, Internet of Things (IoT), WebAssembly, WebAssembly System Interface (WASI)
 I Introduction
@@
 A large amount of IoT devices are integral to critical infrastructure [3] and operate on the edge of the internet, where the management of their software and firmware poses significant challenges due to their inherent complexity [4].
-In accordance with the Cyber Resilience Act [5], manufacturers need to ensure, among various factors, that for the duration of the support period of the product, vulnerabilities are handled effectively and that security updates should be available to users for at least the time the product is expected to be in use.
+In accordance with the Cyber Resilience Act [5], manufacturers must ensure effective vulnerability management and provide security updates for at least the time the product is expected to be in use.
 Moreover, the automotive industry is increasingly adopting the practice of wirelessly distributing software updates to vehicles, known as over-the-air updates, where security is of critical importance [6, 7].
 In parallel, Industrial Internet of Things systems often feature devices with operational lifespans of more than 30 years [8, 9].
-To ensure forward compatibility of these systems, they must be able to integrate with newer hardware components, such as modern and updated sensors and displays, since the original hardware usually already reached end of life when they need to be replaced.
-This highlights the necessity for modularity at the driver level.
-Additionally, these systems are frequently tightly integrated with devices that have limited hardware capabilities, which augments the requirements with the demand for minimal overhead.
-Although cloud-native development practices have introduced benefits such as isolation, componentization, containerization, and modern deployment and orchestration platforms, these advances are difficult to apply to edge environments and even more so for IoT devices, embedded systems, and microcontrollers [10].
-Ensuring drivers are up-to-date and secure is crucial for supply chain security, as they frequently have elevated system privileges or direct access to hardware.
-W3C [11] WebAssembly (Wasm) [12] and the WebAssembly System Interface (WASI) [13] emerges as a promising solution to shrink the trusted computing base and further enhance embedded development and maintainability by providing a trusted runtime, modern development methodologies, toolchains and software development kits.
+To ensure forward compatibility, these systems must be able to integrate with newer hardware components, such as updated sensors and displays, since the original hardware usually reaches end of life when they need to be replaced.
+This necessitates modularity at the driver level.
+Additionally, these systems are often tightly integrated with devices that have limited hardware capabilities, increasing the need for minimal overhead.
+Although cloud-native development practices have introduced benefits such as isolation, componentization, containerization, and orchestration platforms, these advances are difficult to apply to edge environments and even more so for IoT devices, embedded systems, and microcontrollers [10].
+Ensuring drivers are secure and up-to-date is crucial for supply chain security, as drivers often have elevated system privileges or direct access to hardware.
+W3C [11] WebAssembly (Wasm) [12] and the WebAssembly System Interface (WASI) [13] emerges as a promising solution to shrink the trusted computing base and enhance embedded development and maintainability through a trusted runtime, modern development methodologies, toolchains and software development kits.
 Wasm was originally created to execute binary code within the browser to improve the performance of web applications [14], and is now actively used outside the browser via WASI [15].
 WASI facilitates operating system communication within Wasm by providing a set of portable Application Programming Interfaces (APIs).
-However, applications on embedded and IoT devices require access to hardware interfaces to enable communication with the system in which they operate.
-Wasm does not currently meet the demands of cyber-physical systems, particularly in accessing connected hardware, and requires standardized WASI interfaces to be implemented in Wasm runtimes to maintain its portability and enable communication with hardware across diverse devices.
+However, embedded and IoT applications require access to hardware interfaces to enable communication with the system in which they operate.
+Wasm does not currently meet the demands of cyber-physical systems, particularly in accessing connected hardware.
+Standardized WASI interfaces are needed in Wasm runtimes to ensure portability and enable communication with hardware across diverse devices.
 In this work, we propose standardized WASI interfaces and proof-of-concept implementations for connecting hardware using I2C and USB, along with security mediation and isolation, on low-powered embedded and IoT devices running a Wasm runtime with pluggable drivers.
-Given that native drivers are inherently insecure due to the fact that they involve raw memory manipulation, our solution aims to reduce the amount of components that need to be trusted, thus enhancing supply chain security.
-Furthermore, our proposal emphasizes the importance of forward compatible devices and modularity at the driver level, ensuring that the system remains adaptable and secure in the face of evolving technological landscapes.
-Specifically, this paper aims to answer the following questions:
+Our solution aims to enhance supply chain security by reducing the number of components that need to be trusted, addressing the inherent insecurity of native drivers.
+Furthermore, we emphasize the importance of forward compatibility and modularity at the driver level to ensure adaptability and security in evolving technological landscapes.
+Specifically, this paper aims to answer the following Research Questions:
 RQ 1: How to enable WebAssembly applications outside the browser to interact with hardware interfaces such as I2C and USB?
-RQ 2: How to apply access control security measures to ensure authorized access to hardware interfaces?
-RQ 3: How to standardize cyber-physical WebAssembly interfaces to ensure future compatibility?
-RQ 4: How does interacting with hardware interfaces through WebAssembly impact performance compared to using native solutions?
-The remainder of this paper is structured as follows: Section II investigates the state of the art concerning the use of WebAssembly on IoT and edge devices.
-Section III details our proof-of-concept architecture for supporting I2C and USB hardware interfaces within WebAssembly while Section IV discusses the standardization efforts for the WASI proposals, highlighting key decisions made during their development.
-Section V illustrates the feasibility of the WASI proposals through the development of multiple proof-of-concept implementations and Section VI presents the evaluation results of these implementations.
-Finally, Section VII concludes the paper.
+RQ 2: How to apply access control security measures in WebAssembly to ensure authorized access to hardware interfaces?
+RQ 3: How does our solution to interact with hardware interfaces through WebAssembly impact performance compared to native solutions?
+We begin by positioning our work within the existing literature in Section II. Then, we provide a threefold contribution to the state of the art. Specifically, our Contributions are:
+Contribution 1: Develop a novel solution to enable WebAssembly-based workloads on IoT and edge devices to interact with (I2C and USB) hardware interfaces (Section III).
+Specifically, our solution tackles the challenge of giving sandboxed WebAssembly applications mediated access to device resources.
+Contribution 2: Demonstrate the feasibility of our solution through the development of multiple proof-of-concept implementations (Section IV).
+Contribution 3: Evaluate our solution by conducting benchmarks and measuring performance overhead compared to native hardware access (Section V).
 II Related Work
@@
 Singh et al.
-[16] developed the WARDuino virtual machine (VM), which exposes Arduino APIs for GPIO (General-Purpose Input/Output), SPI (Serial Peripheral Interface), ADC/DAC (Analog to Digital and Digital to Analog Converter), and PWM (Pulse Width Modulation) to the Wasm runtime.
+[16] developed the WARDuino virtual machine (VM), which exposes common IoT Arduino APIs to the Wasm runtime.
 WARDuino only facilitates the use of Wasm in, as its name suggests, Arduino-based projects.
@@
 Aerogel enhances the security and provides resource management in multi-tenant Wasm runtime environments, where each Wasm application is considered a tenant.
-Aerogel allows sensor and actuator interaction by exporting high-level sensing and actuation functions to the WebAssembly Micro Runtime (WAMR) [18].
-A user-defined access control specification sheet allows one to set restrictions on specific I/Os, limitations on the number of accesses within a specified time frame, an upper bound on memory and CPU usage, and the maximum number of applications having shared access to a specific hardware component.
+A user-defined access control specification sheet allows one to set restrictions on specific inputs/outputs (I/Os), limitations on the number of accesses within a specified time frame, an upper bound on memory and CPU usage, and the maximum number of applications having shared access to a specific hardware component.
 Access to the requested resource will only be granted if all access control conditions are met.
-WiProg [19] is a framework designed to simplify integrated IoT programming and application development for sensor-to-edge systems by employing Wasm, with a focus on application offloading.
-It allows for a monolithic application development approach which allows WiProg to perform offloading of the application based on compiler annotations and offloading policies.
-WiProg provides hardware-agnostic APIs to access peripherals on IoT devices, but these are limited to analog I/O, digital I/O, and Universal Asynchronous Receiver-Transmitter (UART) APIs, thus lacking support for other important protocols like I2C.
+WiProg [18] is a framework that simplifies integrated IoT programming and application development for sensor-to-edge systems using Wasm, focusing on application offloading.
+It allows for a monolithic application development approach, enabling offloading based on compiler annotations and policies.
+WiProg provides hardware-agnostic APIs for analog I/O, digital I/O, and Universal Asynchronous Receiver-Transmitter (UART), but lacks support for important protocols like I2C.
 Li et al.
-[20] proposed WAIT, a lightweight Wasm runtime for resource-constrained IoT devices for device-cloud integrated applications.
-It enables Wasm ahead-of-time (AOT) compilation targeting resource-constrained devices.
-When a device running WAIT receives a Wasm module binary, it performs on-device AOT compilation, checks sandboxing guarantees, and finally executes the module.
-To cope with the limited resources, the compile-time memory footprint is reduced by streamed look-back compilation.
-To access peripherals and I/O, native APIs are provided by the WAIT runtime and are exposed via the import mechanism of Wasm.
-The AOT compiler then performs binary rewriting techniques to replace the instructions that call the imported function with assembly code that directly accesses the peripheral, reducing the overhead associated with calling an external function.
-Finally, Wasmachine [21] is a bare-metal operating system designed for embedded devices.
-It operates without protection rings by running everything in kernel mode, allowing for zero-cost system calls.
-The Wasmachine kernel exports system calls in WASI prototypes and runs each Wasm application as a kernel thread, allowing the application to invoke system calls as normal function calls without incurring additional context switching costs.
-Wasmachine allows access to hardware by using WASI, where access permissions are configured on a per-application basis and none are granted by default.
-Despite its innovative approach, the system has certain limitations, including limited support of CPU architectures and the reliance on a Unix-like monolithic kernel architecture, which can limit the feasibility on certain IoT or edge devices.
+[19] proposed WAIT, a lightweight Wasm runtime for IoT devices in device-cloud integrated applications.
+It enables Wasm ahead-of-time (AOT) compilation on resource-constrained devices.
+Upon receiving a Wasm module binary, WAIT performs on-device AOT compilation, checks sandboxing guarantees, and executes the module.
+Native APIs for peripherals and I/O are exposed via the Wasm import mechanism.
+The AOT compiler then uses binary rewriting to replace imported function calls with direct assembly code, reducing external function call overhead.
+Finally, Wasmachine [20] is a bare-metal operating system for embedded devices, running everything in kernel mode for zero-cost system calls.
+It exports system calls in WASI prototypes and runs each Wasm application as a kernel thread, allowing the application to invoke system calls as normal function calls without additional context switching costs.
+Hardware access is provided using WASI with per-application permissions.
+Despite its innovative approach, it has certain limitations, including limited CPU architecture support and the reliance on a Unix-like monolithic kernel architecture, which may affect its feasibility on some IoT or edge devices.
 III Architecture
-Fig.
-1 shows a diagram of the proof-of-concept architecture designed by the authors to enable I2C and USB hardware interface support in Wasm applications.
+The proof-of-concept architecture designed by the authors to enable mediated (I2C and USB) hardware interface support in Wasm applications is shown in Fig.
+1.
 Both the application and the peripheral device drivers run as WebAssembly components.
@@
 The host component itself communicates with the host operating system (OS) and uses OS-specific APIs for hardware access.
-The remainder of this section examines important concepts in more detail.
+The remainder of this section examines important concepts of our novel solution in more detail.
 Figure 1: Schematic representation of the proof-of-concept architecture that enables I2C and USB hardware interface support in Wasm applications.
@@
 For example, the USB world contains a WIT description of all the system calls that Wasm components can use to communicate with USB devices.
-This world is a high-level interface contract between the runtime and the guest component and can be used by tools such as wit-bindgen [22] to automatically generate bindings for calling these APIs from various languages.
-The source code of the guest components uses these generated bindings to implement the world and to call the imported APIs, allowing them to interact with hardware devices supported by the interface.
+This world is a high-level interface contract between the runtime and the guest component and can be used by tools such as wit-bindgen [21] to automatically generate bindings for calling these APIs from various programming languages.
+The source code of the guest components uses these bindings to implement the world and to call the imported APIs, allowing them to interact with hardware devices supported by the interface.
 The runtime host component can use these bindings to implement and export the functions that make up the API.
@@
 The host component is the host-side implementation of the system call API.
-The host component is embedded in the runtime and runs as native code directly on the underlying machine, as it requires access to native APIs of the host operating system to communicate with hardware devices.
+It is embedded in the runtime and runs as native code directly on the underlying OS, as it requires access to native APIs of the host OS to communicate with hardware devices.
 As such, the host component is highly dependent on the specific host OS interfaces.
@@
 Any new system or runtime that wishes to support certain worlds, will require a modified implementation of the host component.
-However, this burden can be lessened by using higher-level cross-platform libraries to implement the host component, rather than depending directly on the operating system through system calls.
+However, this burden can be lessened by using higher-level cross-platform libraries to implement the host component, rather than depending directly on the OS through system calls.
 This way, one implementation of the host component can be compiled for multiple operating systems.
-Depending on the specific implementation, it may also be possible to target multiple processor architectures running one of the supported operating systems using a single implementation.
-Unfortunately, using a higher-level library is not always feasible, especially on certain low-powered embedded devices due to device-specific (operating system) limitations, requiring a custom host component implementation for those devices.
+Unfortunately, using a higher-level library is not always feasible, especially on certain low-powered embedded devices due to device-specific limitations, requiring a custom host component implementation for those devices.
 III-C Guest Components
@@
 III-D Runtime
-The Wasm runtime resolves the imports and exports for all the components and links them together.
+The Wasm runtime resolves the imports and exports for all components and links them together.
 To successfully resolve imports and exports, the runtime must support the WIT world that a component implements.
-For example, this can be achieved by having a host component that provides implementations for the imported functions required by the guest component.
+This can be achieved by a host component providing the necessary implementations for the guest component’s imported functions.
 After linking, the guest component has access to the imported functions from the world it implements.
-The runtime also has access to the functions that the world exports.
+The runtime, in turn, has access to the functions that the world exports.
 To begin execution, the runtime must call one of those exported functions.
-Since interface worlds are not concerned with how a component starts execution, this is delegated to another world.
-More specifically, the runtime supports the wasi-cli [23] world to specify the entry-point of the guest component.
-Popular Wasm runtimes include Wasmtime [24] and WAMR [18], which are both being developed by the Bytecode Alliance [25].
+Since interface worlds are not concerned with how a component starts execution, this is delegated to the wasi-cli [22] world to specify the entry-point.
+Popular Wasm runtimes include Wasmtime [23] and WebAssembly Micro Runtime (WAMR) [24], which are both being developed by the Bytecode Alliance [25].
 III-E Capability-Based Security
@@
 If a resource is not provided, the module cannot interact with it, implicitly revoking the capability for that resource.
-For example, WASI file system APIs apply capability-based security so that applications can only access files and directories to which they have been granted access [26].
-Otherwise, the resource does not exist.
 Although WASI enables the implementation of capability-based security for system interfaces, it does not stipulate the level of granularity at which these security controls should be constructed.
@@
   • Device level: Allow only devices with specific vendor and product ID pairs.
-  • Configuration level: Allow only devices with specific configuration types. For example, allow only self-powered devices.
-  • Interface level: Allow only interfaces with specific property types. For example, allow only interfaces that have the mass storage class code.
-  • Endpoint level: Allow only endpoints with specific property types. For example, allow only endpoints with an IN direction that restricts the guest to receiving data only.
+  • Configuration level: Allow only devices with specific configuration types, e.g., self-powered devices.
+  • Interface level: Allow only interfaces with specific property types, e.g., those with the mass storage class code.
+  • Endpoint level: Allow only endpoints with specific property types, e.g., those with an IN direction that restricts the guest to receiving data only.
 The capability-based security model is inherently defined by the API, which also determines the granularity of the security measures.
-This implies that any access control measures must define the resources to be protected, which is impractical for fine granularity and sometimes impossible when control layers are orthogonal or complementary.
-Consequently, we propose an extension of capability-based security with explicit access control list security.
+This implies that access control must define the resources to be protected, which is impractical for fine granularity and challenging when control layers are orthogonal or complementary.
+Consequently, we propose an extension of capability-based security with explicit access control lists.
 III-F ACL-Based Security
-Access control lists provide an additional layer of granular and orthogonal access control, complementing capability-based access control.
+Access control lists add a layer of granular and orthogonal access control, complementing capability-based access control.
 Unlike capability-based access control, which requires changes to the API, ACLs can be implemented without affecting the existing API.
+For example, an ACL rule might allow an application to read from a device but not write to it.
 Furthermore, ACLs offer greater flexibility and can be fine-tuned to the environment and runtime as needed.
-Based on these advantages, we propose that standardized interfaces use capability-based security for high-level concepts where the benefits of access control are clear and use ACL-based access control to further restrict permissions.
-The implementation of ACL-based access control necessitates the incorporation of logic specific to the runtime, environment, and peripheral.
+Based on these advantages, we propose that standardized interfaces use capability-based security for high-level concepts and use an additional ACL to further restrict permissions.
+Our implementation of ACL-based access control necessitates the incorporation of logic specific to the runtime, environment, and peripheral.
 Whenever the runtime creates a new module instance, an I/O context is created and linked to the corresponding execution environment.
-Whenever a device interaction is performed, the corresponding I/O context is retrieved from the execution environment.
-Depending on the type of device interaction and the I/O context’s properties, the action is permitted or denied.
+Whenever a device interaction is performed, the corresponding I/O context is retrieved from the execution environment, and depending on the type of device interaction and the I/O context’s properties, the action is permitted or denied.
 The I/O context’s properties may refer to permitted interaction types, such as read or write.
@@
 Furthermore, the context may specify how connections should be established with end devices, preventing communication with other peripherals sharing the same bus.
-IV System Interface Standards
-In order to ensure compatibility with future toolchains and platforms, it is important to standardize the cyber-physical WebAssembly interfaces.
-Standardization will enable the integration of support for these interfaces into runtimes, compiler toolchains, HAL libraries, and drivers.
-As part of this research, the authors joined the WebAssembly System Interface Subgroup of the W3C WebAssembly Community Group [27] to initiate the standardization process [^1] https://github.com/WebAssembly/meetings/blob/main/process/phases.md for these interfaces.
-These efforts focused on two interfaces commonly used in IoT and embedded systems: I2C and USB.
-At the time of writing, the WASI-I2C [28] proposal is in phase 2 and the WASI-USB [29] proposal is in phase 1 of the standardization process.
-In the following, we outline both proposals, emphasizing certain key decisions that were made during the development of the proposals.
-IV-A WASI-I2C
-The WASI-I2C proposal is co-championed by representatives from both imec and Siemens, and is currently in the second phase of the standardization process, with ongoing efforts to fulfill the criteria to pass the vote to the third phase.
-The WASI-I2C proposal contains three WIT files adhering to the component model: delay, i2c, and world.
-The delay and i2c files are closely aligned with the interfaces from embedded-hal [30] v1.0, which served as inspiration, although the use of the embedded-hal Rust crate is not required to implement the proposal.
-The world file provides a way to import all the described interfaces of the proposal at once, thus currently importing all interfaces from the delay and i2c files.
-The proposal describes two handles that dynamically identify and provide access to resources, similar to a file descriptor.
-One handle is designated for I2C and the other for delays.
-These handles offer extensive access to the resource, meaning you currently cannot restrict the I2C connection to specific addresses.
-While a more fine-graded control model is feasible, consultations with the embedded special interest group (SIG-Embedded) [31] and industry partners have indicated that there is currently no demand for this.
-Because there exist various ways to obtain an I2C handle, the i2c interface lacks an explicit constructor for the i2c resource.
-Instead, worlds that utilize this interface should provide their own mechanism to obtain I2C handles, typically by passing handles into exported functions as parameters or by having imported functions return them.
-The i2c resource includes essential methods such as read and write to interact with the I2C device.
-Although delaying program execution is not inherently associated with I2C communication, certain guest implementations depend on it, which is why it is also included in the current proposal.
-In the future, delays could either be separated into their own proposal or merged with other embedded-related proposals to avoid the need for each delay dependent proposal to establish its own standards.
-The main downside of merging proposals is that it could significantly increase the workload for the proposal champions, who would need to be interested and committed to maintaining these additional communication protocols.
-However, having centralized proposals could streamline the process for potential implementers.
-A key element of a proposal is its portability criteria.
-These criteria demonstrate that the champions’ specification is not overly tailored to a specific platform.
-Typically, there need to be at least two entirely independent implementations in separate Wasm runtimes.
-However, given the diversity of platforms that can interface with I2C, this approach is not feasible for the WASI-I2C proposal.
-The champions, along with representatives from Siemens, Intel, Aptiv, Xiaomi, Amazon, Midokura, Sony, Atym, and Bosch, have collectively established a set of portability criteria for the proposal, as outlined in Table I.
-To minimize the potential variety in memory size, flash size, and processor speed, reference hardware was provided for each criterion.
-Although the Intel x86 architecture is not explicitly mentioned, its support is implied since Wasmtime is being developed using this architecture.
-The NUCLEO-F412ZG was selected as it serves as Siemens’ reference board.
-The remaining two targets were selected to demonstrate compatibility with both a real-time operating system (RTOS) and Linux, as well as ARM and RISC-V architectures.
-In addition to the target requirements, it is crucial that the interface is designed in a way that minimizes its memory usage, ensuring that there is sufficient memory available for the application itself.
-TABLE I: WASI-I2C Proposal Portability Criteria
-Platform Architecture Reference Hardware
-Linux ARM64 Raspberry Pi 3B
-RTOS RISC-V ESP32-C2
-RTOS ARM32 NUCLEO-F412ZG
-IV-B WASI-USB
-The WASI-USB proposal is championed by imec and is currently in the first phase of the standardization process.
-The WASI-USB proposal contains a preliminary version of the USB API interface and is described using the WIT format. The key components of the proposal are outlined below.
-The usb interface contains a usb-device and a device-handle resource.
-The static enumerate function of the usb-device resource allows the caller to retrieve all available USB devices on the bus, returning them as a list of usb-device instances.
-Moreover, there are three other methods that can be used on any of these instances.
-The device-descriptor and configurations methods provide additional information about a device and are always available without requiring exclusive access to the device.
-Conversely, the open method is only available under specific conditions.
-This method establishes a context for device communication.
-It is important to note that opening a device can fail for various reasons.
-For instance, some operating systems may refuse to open a device if the executable lacks the necessary permissions.
-These permissions are managed at the Wasm runtime level, rather than being specific to individual Wasm modules.
-To develop a robust and user-friendly API, functions that require an ‘open’ device are scoped to the device-handle resource, which is only obtainable when the device is opened successfully.
-The device-handle resource encompasses all the methods necessary for device communication, with most methods closely related to similar libusb [32] functions.
-The events interface allows a guest to poll for connection event updates.
-This allows the guest to detect connect and disconnect events and react accordingly.
-The events interface includes a single update function that returns a device-connection-event variant.
-The returned case of device-connection-event will vary based on the specific event that has occurred, for example pending when no new event update is available.
-With the anticipated async support in WASI Preview 3 [33, 34], the update function can be adapted to receive event updates through an asynchronous stream rather than polling.
-IV-B1 Security Model Considerations
-The current WASI-USB proposal applies the capability-based security model exclusively at the device level.
-In this approach, only devices with specified vendor and product ID pairs are permitted.
-This is the most straightforward option and is similar to how WebUSB protects devices [35].
-Community feedback will be essential to determine whether the capability-based security model should be extended to other levels, based on its perceived usefulness.
-IV-B2 Exclusive Device Access
-Multiple components may attempt to access the same USB device simultaneously.
-However, this should be avoided, as each component typically expects exclusive access to a device.
-To address this, the runtime should monitor which components are using specific devices and prevent other components from accessing a device that is already in use.
-This can be achieved by returning an error when a component attempts to open a device that is already in use.
-This approach is consistent with most operating systems, where only one program can access a USB device at any given time.
-V Implementation Details
-To demonstrate the feasibility of the proposed cyber-physical WebAssembly architecture, we developed several proof-of-concept implementations, each addressing different aspects of the proposed solution.
-Our primary objectives were to validate the proposed approach and identify potential opportunities or shortcomings in the underlying architecture and standard methodologies.
-In particular, this research presents three distinct proof-of-concept implementations.
-  • I2C in Wasmtime: Demonstrates driving an I2C seven-segment display from WebAssembly running on a Linux host and on a Raspberry Pi Pico microcontroller, as well as reading temperature and humidity data from an HTS221 sensor on a Raspberry Pi 3B.
-  • I2C in WAMR: Provides a proof of concept that can read data from an HTS221 sensor.
-Note that this solution does not yet use the component model or capability-based security due to the current limitations of WAMR, including the lack of support for WASI Preview 2.
-Instead, an ACL-based Security System was implemented.
-  • USB in Wasmtime: Showcases the communication of WebAssembly applications with USB devices, including a USB flash drive using a WebAssembly-based driver and an Xbox controller using a separate driver.
-The remainder of this section describes each proof-of-concept implementation.
-V-A I2C in Wasmtime
+IV Implementation Details
+To validate our proposed cyber-physical WebAssembly architecture, we developed three proof-of-concept implementations, each addressing different aspects of the solution.
+Our primary objectives were to confirm the feasibility of the approach and identify potential opportunities or shortcomings in the underlying architecture and standardization methodologies.
+  • I2C in Wasmtime: Demonstrates driving an I2C seven-segment display from Wasm on a Linux host and a Raspberry Pi Pico, and reading temperature and humidity data from an HTS221 sensor on a Raspberry Pi 3B.
+  • I2C in WAMR: Demonstrates reading data from an HTS221 sensor.
+Due to WAMR’s current limitations, including the lack of WASI Preview 2 support, this solution uses an ACL-based security system instead of the component model or capability-based security.
+  • USB in Wasmtime: Demonstrates Wasm applications communicating with USB devices, including a flash drive and an Xbox controller using Wasm-based drivers.
+IV-A I2C in Wasmtime
 The I2C in Wasmtime proof-of-concept implementation [26] is made up of several components: the host implementation, two WIT worlds and two guest implementations.
-The host implementation handles the actual I2C connection with the device, whereas the guest implementations are responsible for device-specific logic, such as reading the temperature or writing the time to the display.
-The host and guest components interact with each other through two WIT worlds, one for the seven-segment display and the other for the HTS221 sensor.
-These WIT worlds import the necessary interfaces from the WASI-I2C proposal that are implemented by the host, allowing the guest to communicate with the I2C device.
-Additionally, these WIT worlds export device-specific functions, such as get-temperature, that are implemented and made available by the guest.
+The host implementation handles the I2C connection with the device, whereas the guests are responsible for device-specific logic, such as reading temperature or displaying time.
+The host and guest components interact through two WIT worlds, one for the display and the other for the HTS221 sensor.
+These WIT worlds import the necessary interfaces from the WASI-I2C [27] proposal that are implemented by the host, allowing the guest to communicate with the I2C device.
+Additionally, the WIT worlds export device-specific functions implemented by the guest.
 Although these worlds were written for our specific proof of concepts, they aim to be as generic as possible to allow future support for other displays and sensors.
-For example, the display’s write function can accept any string, even though our display only has four seven-segment characters.
-The guest implementations do not make use of Rust’s std library because they only need specific types that can be imported individually, which also helps minimize the size of the compiled Wasm binaries.
-The guest implementation for the seven-segment display directly writes the necessary bytes to control it, while the guest component for the HTS221 sensor leverages the embedded-hal compatible hts221 [37] Rust crate for simplicity.
+The guest implementations do not use Rust’s std library, only importing necessary types individually, which helps minimize the size of the compiled Wasm binaries.
+The guest component for the seven-segment display directly writes bytes to control it, while the HTS221 sensor component leverages the embedded-hal [28] compatible hts221 [29] Rust crate for simplicity.
 To adapt the embedded-hal API to the slightly different generated bindings, the wasi-embedded-hal [30] crate was developed.
-V-B I2C in WAMR
+IV-B I2C in WAMR
 WAMR currently does not support WASI Preview 2.
-Consequently, the component model and WIT cannot be used in WAMR, which means that a standardized I2C connection cannot be passed to the guest because WIT resources are not available.
-As a workaround, the I2C connection is kept global on the host side.
-For the host to be able to bind with the guest, it has to register the function that the guest imports and find the functions that the guest exports.
-Parameters are passed to the guest’s exported functions using a vector of WasmValue.
-Internally, a WasmValue is represented as an array of 32-bit unsigned integers.
-The ACL security system for the I2C implementation employs a struct i2c_context to store access permissions.
-A reference is registered in the user_data field of the execution environment, which contains the device address, indicating that the specific module is permitted to communicate with the device.
-As it is necessary to reestablish the I2C connection on each I2C API call, a check for permissions can be performed.
-Given the frequency with which I2C functions are called in embedded applications, it is necessary to ensure that the lookup time is constant.
-To this end, the allowed addresses are stored in two 64-bit fields spanning the full range of 128 valid I2C addresses, thereby ensuring that the lookup time remains consistent.
-V-C USB in Wasmtime
-The proof-of-concept implementation [39] of USB in Wasmtime includes a host implementation along with several guest implementations, such as interfacing with a USB flash drive or an Xbox controller.
-The main concept of the USB implementation in Wasmtime mirrors that of the I2C implementation in Wasmtime: the host component is platform dependent and interacts with the underlying system to access the USB resource, while the guest components handle device-specific logic, such as reading the joystick positions of an Xbox controller.
+Consequently, the component model and WIT are not available, which means that a standardized I2C connection cannot be passed to the guest.
+As a workaround, the I2C connection remains global on the host side.
+For the host to bind with the guest, it has to register the function that the guest imports and find the functions that the guest exports.
+IV-C USB in Wasmtime
+The proof-of-concept implementation [31, 32] of USB in Wasmtime includes a host implementation along with several guest implementations.
+The main concept of the USB implementation in Wasmtime mirrors that of the I2C implementation in Wasmtime: the host component is platform dependent and interacts with the underlying system to access the USB resource, while the guest components handle device-specific logic, such as reading the inputs of an Xbox controller.
 To facilitate the development of the host component and enhance its platform independence, rusb [33], a Rust wrapper around libusb [34], is used.
-The USB flash drive guest component implements a range of commands for USB Attached SCSI devices, including basic file system operations like listing directory contents and reading files.
-To streamline the proof-of-concept implementation, it utilizes the fatfs [41] Rust crate for file system support.
-The file system support is fully integrated within the guest component itself, without relying on the host operating system’s file system capabilities.
+The USB flash drive guest component supports various USB Attached SCSI device commands, including basic file system operations like listing directory contents and reading files.
+The file system support is fully integrated within the guest component itself using the fatfs [35] Rust crate, without relying on the host operating system’s file system capabilities.
 The Xbox guest component continuously monitors the current state of the controller through its USB interrupt pipe and writes button presses and joystick positions to standard output.
-VI Evaluation Results
+V Evaluation Results
 This section assesses the practicality of our proposals in real-world scenarios.
@@
 Subsequently, various benchmarks are performed to identify any performance implications.
-VI-A I2C Performance Evaluation
+V-A I2C Performance Evaluation
 The I2C proof of concept for the seven-segment display is evaluated on a Raspberry Pi 4B, whereas the HTS221 sensor proof of concept is evaluated on a Raspberry Pi 3B.
 The display evaluation will compare the native, Wasmtime, and WAMR implementations, while the HTS221 sensor evaluation will compare the native and Wasmtime implementations.
-All benchmark results are available on GitHub [36].
+All benchmark results are available [26] on GitHub.
 To assess performance overhead, we conduct 100 iterations of writing a four-digit number to the display and also reading the temperature from the HTS221 sensor.
@@
 In contrast, the WAMR implementation incurs only a minimal overhead of 50 $\mu$s (+1.7%).
-Most of the additional time consumed by Wasmtime can be attributed to functions associated with Cranelift [42], a compiler backend utilized by Wasmtime for both just-in-time and ahead-of-time (AOT) compilation.
+Most of the additional time consumed by Wasmtime can be attributed to functions associated with Cranelift [36], a compiler backend utilized by Wasmtime for both just-in-time and ahead-of-time compilation.
 Figure 2: Mean cold start execution times for writing a four-digit number to the display and reading the temperature from the HTS221 sensor, comparing native, Wasmtime, and WAMR implementations.
-VI-B USB Performance Evaluation
+V-B USB Performance Evaluation
 The USB proof of concept is evaluated on three systems: one with an AMD Ryzen 7 5800H and USB3.0 running x86 Linux 6.8.9, a Raspberry Pi 3B+ with USB2.0 running AArch64 Linux 6.8.0, and a third with an AMD Ryzen 7 5800X3D and USB3.0 running x86 Windows 10 22H2.
-All benchmark results are available on GitHub [43].
+All benchmark results are available [37] on GitHub.
 To assess the runtime size and memory usage overhead introduced by the USB host component, we compare the Wasmtime runtime both with and without the USB host component.
 Our measurements show that the runtime size overhead is minimal, adding only 160 KiB (+1.1%) for the x86 Linux runtime, 220 KiB (+1.7%) for the AArch64 Linux runtime, and 275 KiB (+1.9%) for the Windows runtime.
-The memory usage overhead is also minimal for the Linux runtimes, comprising only 459 KiB (+1.7%) for the x86 Linux runtime and 295 KiB (+1.5%) for the AArch64 Linux runtime.
-The Windows version of the runtime incurs a more severe memory overhead of 1427 KiB (+11.0%), which can be attributed to an entirely different backend being used by libusb on Windows.
+Memory usage overhead is also minimal for the Linux runtimes, comprising only 459 KiB (+1.7%) for the x86 Linux runtime and 295 KiB (+1.5%) for the AArch64 Linux runtime.
+The Windows runtime incurs a more severe memory overhead of 1427 KiB (+11.0%), which can be attributed to an entirely different backend being used by libusb on Windows.
 The impact on throughput is assessed using a USB flash drive to read or write as many data blocks as quickly as possible.
@@
 Some outliers have been excluded to facilitate the comparison of the median RTTs.
-The overhead introduced by WebAssembly is relatively modest.
+The overhead introduced by Wasm is relatively modest.
 On x86 Linux, it results in a median RTT that is only 7 $\mu$s (+3.5%) longer than the native implementation.
@@
 The reported times represent the median round-trip time.
-VII Conclusion
+VI Conclusion
 With the growing popularity of WebAssembly on embedded and IoT devices, the need for standardized hardware interface support in WebAssembly arose.
-This paper introduces two WebAssembly System Interface proposals, WASI-I2C and WASI-USB, enabling the use of I2C and USB hardware interfaces within WebAssembly.
-The cyber-physical WebAssembly proof-of-concept architecture is described in depth, including a solution for authorized hardware interface access.
+This paper introduces two WebAssembly System Interface proposals, WASI-I2C and WASI-USB, enabling the use of I2C and USB within WebAssembly.
+Our cyber-physical WebAssembly proof-of-concept architecture is described in depth, including a solution for authorized hardware interface access.
 Hardware interface access is managed through the capability-based security mechanism of WASI, and is extended with explicit ACL-based security for more granular access control.
-To ensure our proposals remain compatible with future toolchains and platforms, we have initiated the standardization process in collaboration with industry partners.
-Development details for the standardized proposals are provided, highlighting key decisions made throughout the process.
 Proof-of-concept implementations showcase the feasibility of the proposed architecture and are evaluated to confirm that they meet functional requirements and to identify any performance impacts.
@@
 Acknowledgment
-We extend our heartfelt gratitude to everyone who contributed to these proposals: Valentin Olpp, Dan Gohman, Emiel Van Severen and the Bytecode Alliance & WASI subgroup of the W3C WebAssembly Community Group.
+We extend our heartfelt gratitude to everyone who contributed to these standardization proposals: Valentin Olpp, Dan Gohman, Emiel Van Severen and the Bytecode Alliance & WASI subgroup of the W3C WebAssembly Community Group.
 This work received funding from the Smart Networks and Services Joint Undertaking (SNS JU) under the Horizon Europe research and innovation program of the European Union under Grant Agreement number 101139067 (https://elasticproject.eu/).